The demand equation for a certain type of printer is given by D= -200p + 35,000
The supply equation is predicted to be S= -p^2+400p-20,000
Find the equilibrium price.

Answer by stanbon(48574)

The demand equation for a certain type of printer is given by D= -200p + 35,000
The supply equation is predicted to be S= -p^2+400p-20,000
Find the equilibrium price.
-------------
Equilibrium means demand = supply.
EQUATION:
-200p + 35,000 = -p^2+400p-20,000
p^2-600p+55000 = 0
p = [600+sqrt(600^2-4*1*55000]/2
Positive answer:
p = $487.08
=============
Cheers,
Stan H.